OpenStudy (anonymous):

Help please! Finals tomorrow! I need help with factoring. x^3-x^2-9x+18

OpenStudy (dan815):

you gotta take a guess, like guess an x that might work good guesses can be made by look at the coefficients of x^3 and the constant term

OpenStudy (dan815):

you guess +/- the factors of the coeffecient on X^3 or the constant in the end

OpenStudy (dan815):

so like +/-1 or +/-2 or +/-3 or +/-6 and so on one of those will work for sure if they expect this to be factored

OpenStudy (dan815):

guess that and then say +3 worked then you would do synthetic division with (x+3)

OpenStudy (dan815):

this does not have good factors wuut

OpenStudy (dan815):

you are going to have to use stuff like newtons method to get a real close approximation

hero (hero):

As far as factoring goes, the best you can do is this: \(x^3 - x^2 - 9x + 18 = x^2(x - 1) - 9(x - 2)\) \(= x^2(x - 1) - 9(x - 1) + 9\) \(= (x - 1)(x^2 - 9) + 9\) \(= (x - 1)(x + 3)(x - 3) + 9\) In the end, you'll have 9 left over no matter what you do